# A general approach to modeling environmental mixtures with multivariate outcomes

### v1

@@ 2.2 Prior distributions inducing clustering
Here $\pi_{a}^{\beta}$ and $\pi_{b}^{\theta}$ are assigned the same stick-breaking prior distributions as in the independent DP formulation described above, and, as before, the concentration parameters $\alpha_{\beta}$ and $\alpha_{\theta}$ control the level of clustering for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. However, $\rho$ now quantifies the degree of dependence in clustering between $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. When $\rho=0$, the model reduces to the independent DP prior case. When $\rho\to\infty$, $P(Z_{\beta,j}=a,Z_{\theta,j}=b)=1$ when $a=b$ and 0 otherwise; that is, either both $\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}}$, or $\bm{\beta}_{kj}\neq\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}\neq\bm{\theta}_{k^{\prime}j^{\prime}}$.
We specify the centering distributions $G_{0,\beta}$ and $G_{0,\theta}$ to be the smoothness inducing priors detailed in Section 2.1. To complete the model specification, we adopt default gamma priors on ${\alpha_{\beta},\alpha_{\theta}}$, $\rho$, and smoothness parameters ${\lambda^{\beta},\lambda^{\theta}}$. For full details, see Appendix A.
3 Extension to Multiple Index Models
When exposures are not time varying, the proposed approach extends immediately to a (multivariate) single index model. Let ${\mathbf{x}}_{i}$ now represent an $L$-vector of different exposures measured for the $i^{th}$ observation. (Note: we represent the number of exposures by $L$ here rather than $P$ as in the previous section.) The multivariate single index model is:
@@ 3 Extension to Multiple Index Models
for $k=1,\dots,K$. In this model, ${\mathbf{x}}_{i}^{T}{\bm{\omega}}_{k}$ represents a multipollutant index, $f_{k}(\cdot)$ characterizes the shape of the association between the $k^{th}$ outcome and the multipollutant index, and ${{\omega}}_{kj}$ quantifies the relative strength of association between the $k^{th}$ outcome and $j^{th}$ exposure. As before, we may borrow strength across different outcomes by encouraging similarity both in ${\bm{\omega}}_{k}$ and $f_{k}(\cdot)$ via the cluster inducing priors of Section 2.2. Inference follows as above, but without a penalty to encourage smoothness across weights ${\omega}_{kj}$, $j=1,\dots,L$.
This can be overly restrictive, requiring that the association between each exposure and the $k^{th}$ outcome have the same shape, differing only in magnitude. Alternatively, multiple index models (MIMs) have been used for environmental mixtures by partitioning the exposure vector ${\mathbf{x}}_{i}$ (McGee et al.,, 2021), but this requires knowing the true index structure a priori. Instead, we propose to extend this strategy to a (multivariate) MIM of maximum order $P$:
$y_{ik}$ $=\sum_{j=1}^{P}f_{kj}({\mathbf{x}}_{i}^{T}{\bm{\omega}}_{kj})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~{}~{}k=1,\dots,K.$ (2)
This is nearly identical to model (1), but instead of summing over exposures, the summation is over over multipollutant indices that share common exposures $\mathbf{x}_{i}$. Applying the priors proposed in Section 2.2 now encourages clustering across outcomes and indices rather than exposures. This has two key consequences. First, it again improves efficiency by borrowing strength across outcomes. Second, clustering across indices facilitates collapsing to a lower order ($<P$) index model. For example if $P$=2, then when $Z^{\theta}_{k1}=Z^{\theta}_{k2}$ and $Z^{\beta}_{k1}=Z^{\beta}_{k2}$, model (2) reduces to a single index model—without assuming this a priori. This marks a major advantage over standard MIMs, as we need only specify the maximum order $P$, rather than the true index structure (i.e., the number of indices and their composition).
@@ 3.1 Exposure Importance Measures
In the time-varying exposure setting of Section 2, weights $\bm{\omega}_{kj}$ have a natural interpretation as the strength of association at different lags. In a SIM, the weights have the interpretation of being the strength of association between a particular exposure and outcome. In a MIM with unknown index structure, however, $\bm{\omega}_{kj}$ does not have a natural interpretation, and it is difficult to infer which exposures are driving the mixture association. One option is to investigate component-wise exposure response curves when holding all other exposures fixed, but this can be misleading when exposures are highly correlated. To facilitate interpretation, we propose alternative exposure importance measures that can be used for complex mixture models like the proposed MIM.
Rewriting $f^{*}_{k}(\bm{x})=\sum_{j=1}^{P}f_{kj}({\mathbf{x}}^{T}{\bm{\omega}}_{kj})$ in (2) and omitting the subscript $i$ for simplicity, we first define the overall signal-to-noise ratio for the $k^{th}$ outcome as $\text{SNR}_{k}=\frac{Var[f^{*}_{k}(\bm{X})]}{Var(\xi\sigma_{k}u+\epsilon_{k})},$ where the numerator represents the variability explained by the mixture (the signal), and the denominator represents unexplained variability (the noise). This quantifies the overall strength of the association between outcome $k$ and the exposure mixture, but does not provide information about which exposures are driving that association. By further decomposing the numerator, we construct variable importance measures for each exposure-outcome pair:
$\Psi_{kj}=1-\frac{Var_{\bm{X}_{-j}}\bigg{\{}E_{x_{j}\mid\bm{x}_{-j}}\Big{[}f^{*}_{k}(x_{j},\bm{x}_{-j})\Big{]}\bigg{\}}}{Var[f^{*}_{k}(\bm{x})]},$ (3)
@@
where $\bm{x}_{-j}$ represents all exposures except the $j^{th}$ exposure. $\Psi_{kj}$ represents the proportion of the signal explained only by the $j^{th}$ exposure. Naturally, $\Psi_{kj}$ will be small if $x_{j}$ is not predictive of the outcome. But it will also be small if $x_{j}$ is highly collinear with $\bm{x}_{-j}$, as the association with $x_{j}$ can be easily captured by other components. In this way $\Psi_{kj}$ quantifies the importance of the $j^{th}$ exposure to the mixture effect on the $k^{th}$ outcome. In Appendix C we generalize this to arbitrary groups of exposures.
@@
Computing these quantities is straightforward given posterior samples of $f_{k}^{*}(\bm{x})$. Marginal variances are calculated using the empirical distribution of the exposures. Estimating $\Psi_{kj}$ requires integrating over the conditional distribution of $x_{j}$ given the other exposures. If the dimension of $\bm{x}$ is small, we can estimate the conditional mean in (3) non-parametrically:
$\widehat{E}_{x_{j}\mid\bm{x}_{-j}}\{f^{*}_{k}(x_{j},\bm{x}_{-j})\}=\frac{\sum_{i=1}^{n}f^{*}_{k}(x_{ij},\bm{x}_{(-j)})K(\bm{x}_{i(-j)}-\bm{x}_{-j})}{\sum_{i=1}^{n}K(\bm{x}_{i(-j)}-\bm{x}_{-j})}$
@@
where $K(\cdot)$ is an appropriate kernel. Alternatively, when $\bm{x}$ is high-dimensional, one could make parametric assumptions, such as assuming that $x_{j}$ given $\bm{x}_{-j}$ is normally distributed, and the conditional mean can be estimated using any regression approach.
The proposed $\Psi_{kj}$ has connections to a number of related variable importance metrics in the literature. It is closely related to those proposed in Shin et al., (2024), which instead target effect modification and identify variables that modify associations the most. It is also related to metrics used in nonparametric regression such as Shapley values (Shapley,, 1953) or leave-one-covariate-out (LOCO; Lei et al.,, 2018) that target the extent to which mean squared error is reduced when a covariate is omitted. One important distinction is that our proposed metric does not require fitting multiple models (e.g., omitting exposure $j$). Rather, it uses a single fit along with an estimate of the distribution of the exposures. While the distribution of the exposures can be misspecified, we view $\Psi_{kj}$ as a complementary piece of information to aid interpretation, even if the exposure distribution is imperfectly estimated.
4 Posterior sampling of index parameters
We sample from the posterior via MCMC. Using a standard Gibbs sampler, sampling from the full conditionals for most parameters is straightforward (see Appendix B for details). Sampling from the full conditional for the index weights $\bm{\theta}^{*}_{c}$, however, is complicated by its constrained parameter space. We present two approaches to sampling $\bm{\theta}^{*}_{c}$: the first transforms $\bm{\theta}^{*}_{c}$ to an unconstrained space; the second exploits a Taylor approximation to obtain a closed form for the full conditional distribution and ignores the sign constraint.
4.1 Polar Coordinate Reparameterization
The constraints on $\bm{\theta}^{*}_{c}$ makes choosing reasonable proposal distributions challenging in practice. One path forward is to absorb the constraints into the parameterization by transforming to polar coordinates. Following the approach of Park et al., (2005), we let
4.2 Fast Approximate Sampling via Taylor Expansion
The approach of Section 4.1 can be slow to mix in high-dimensional settings. Alternatively, we propose the following approximate method, ignoring the sign constraint and using the Fisher-Bingham distribution on the full hypersphere for $G_{0,\theta}$. A first order Taylor expansion of $f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})$ around the previous value $\tilde{\bm{\theta}}^{*}_{c}$ implies that the sum of squares is approximately
where $\hat{\mathbf{y}}_{k}$ is a vector with $i^{th}$ element equal to
This corresponds to a Fisher Bingham distribution, $\bm{\theta}^{*}_{c}\sim FB({\bm{\eta}}^{T},-\frac{1}{2}\bm{\Lambda}^{-1})$, which can be conveniently sampled from via the simdd package in R (Kent et al.,, 2018). Despite its closed form, directly sampling from a Fisher-Bingham distribution can be computationally challenging in high dimensions. Absent the indicator function, the density in (4) is proportional to a multivariate normal distribution for $\bm{\theta}^{*}_{c}$. As such, in high-dimensional settings we propose to use an approximate update, where we first draw a random variable from a $N\left(\bm{\Lambda}{\bm{\eta}},~{}\bm{\Lambda}\right)$ distribution and then project this value to the unit sphere. Empirically we have found that this approximation leads to very similar samples from the posterior distribution of $\bm{\theta}^{*}_{c}$ and is significantly faster to compute.
@@ 5 Simulation studies
We present simulations demonstrating the effectiveness of the proposed methods in two different contexts. In Simulation A, we consider time-varying exposures and demonstrate the benefit of borrowing strength across exposures and outcomes in DLNMs. In Simulation B, we consider cross-sectional exposure measurements and try to capture complex, non-linear relationships between a mixture of exposures and outcomes using MIMs. In both cases, we average results across 250 simulated datasets.
@@ 5.1 Simulation A: Distributed lag models with multiple exposures
We first focus on settings where exposures are measured repeatedly over time prior to outcomes measured at a single time point. Throughout, we set the number of outcomes to be $K=4$, the number of exposures to be $P=5$, and the number of time points to be $L=52$. We varied the sample size $n=\{500,1500\}$. To induce autocorrelation within each exposure as well as dependence among different exposures, we generated the exposures from a vector autoregressive model. Specifically, we set the exposures at the first time point to be $\bm{X}_{1}=\bm{\epsilon}_{1}$, where $\bm{\epsilon}_{1}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma})$, and then generated subsequent exposures according to:
$\bm{X}_{l}=\bm{M}\bm{X}_{l-1}+\bm{\epsilon}_{l}\ \text{for }l=2,\dots,L,\quad\quad\bm{\epsilon}_{l}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma}).$
We set $\bm{M}=0.85\bm{I}_{p}$ to induce a high degree of correlation over time within exposures, and we set $\bm{\Sigma}_{kj}=0.6^{|j-k|}$ to induce dependence across exposures, as is typical in mixtures of environmental exposures. For a visualization of exposure trajectories, see Appendix D.
@@
Let $\bm{x}_{j}$ represent the vector of $L$ measurements of the $j^{th}$ exposure. Responses were generated from a multivariate normal distribution with mean:
$E[\bm{Y}]$ $=\left[\begin{array}[]{l}f_{1}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{1}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{1}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{3}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{3}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{3}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{4}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{4}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{4}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\end{array}\right],$
@@
where $f_{1}(a)=0.04a$, $f_{2}(a)=0.24(0.3a)^{2}$, $f_{3}(a)=0.09a$, $f_{4}(a)=2\sin(0.2a)$; $\bm{\omega}_{1}$ is flat, $\bm{\omega}_{2}$ is increasing, and $\bm{\omega}_{3}$ is decreasing. The lag profiles were the same across outcomes, and the exposure response functions were the same across exposures. Ultimately, of the $K\times P$=20 different pairs, there were 12 distinct outcome-exposure relationships.
To each simulated dataset we then fit four variations of the proposed approach:
  1. 1.
    The MV-DLNM with dimension reduction as in Section 2.1 (DLNM dimension reduction)
  2. 2.
    The MV-DLNM with no dimension reduction (DLNM)
  3. 3.
    The MV-DLNM without cluster-inducing priors (DLNM no clustering)
  4. 4.
    The DLNM applied to each outcome separately (DLNM separate). This allows for clustering across exposures but not across outcomes.
We report mean squared error (MSE) for the posterior mean of $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$, as well as 90% credible interval coverage, averaged over all parameters and datasets. For the proposed MV-DLNM, we also report pairwise posterior clustering probabilities—$P(\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}}|\bm{y})$ and $P(\bm{\omega}_{kj}=\bm{\omega}_{k^{\prime}j^{\prime}}|\bm{y})$—averaged across datasets.
@@
Figure 1 displays average pairwise clustering probabilities. Overall the proposed framework was able to capture similarity in both $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$. There was a high degree of clustering for the $\bm{\beta}_{kj}$ parameters within outcomes, as characterized by the block diagonal structure in the top left panel, and a low degree of clustering elsewhere. One exception was that there was a moderate degree of clustering between outcomes 1 and 3 when $n=500$; the proposed clustering mechanism was capturing the fact that both had the same shape of exposure response relationship (both linear), despite being of different strengths. This improved with higher sample size, with the proposed approach erroneously clustering outcomes 1 and 3 much less frequently (bottom left panel; Figure 1). Pairwise clustering probabilities were somewhat lower for $\bm{\omega}_{kj}$ (top right panel; Figure 1), as the lag weights were only weakly identified due to high autocorrelation (and exacerbated by correlation among different exposures). Again this issue improved with higher sample size (bottom right panel; Figure 1), with pairwise similarity in lag profiles being captured much more often.

### v2

@@ 2.2 Cluster-Inducing Prior Structure
Here $\pi_{a}^{\beta}$ and $\pi_{b}^{\theta}$ are assigned the same stick-breaking priors as above, and $\alpha_{\beta}$ and $\alpha_{\theta}$ again control the levels of clustering for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. However, $\rho$ now quantifies the degree of dependence in clustering between $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. When $\rho=0$, the model reduces to the independent DP case. When $\rho\to\infty$, $P(Z_{\beta,j}=a,Z_{\theta,j}=b)=1$ when $a=b$ and 0 otherwise; that is, either both $\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}}$, or $\bm{\beta}_{kj}\neq\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}\neq\bm{\theta}_{k^{\prime}j^{\prime}}$.
For $G_{0,\beta}$ we adopt smoothness-inducing priors, as we expect small changes in exposures to lead to small changes in outcomes. We adopt a mean-zero Gaussian with precision matrix $\lambda^{\beta}\bm{\Sigma}^{-1}_{0}$, where $\bm{\Sigma}_{0}$ is a matrix of integrated second derivatives of basis functions to penalize overly rough functions (wood2017generalized), and $\lambda^{\beta}$ is a smoothing parameter. For $G_{0,\theta}$ we adopt ridge penalty priors by default with appropriate constraints (and penalty parameter $\lambda^{\theta}$), but consider an alternative for time-varying exposures in Sec. 3. See Appendix B for details.
2.3 Posterior sampling
We sample from the posterior via MCMC. Using a standard Gibbs sampler, sampling most parameters is straightforward (see Appendix D). Sampling from the full conditional for the index weights $\bm{\theta}^{*}_{c}$, however, is complicated by its constrained parameter space. In Appendix C we present two approaches to sampling $\bm{\theta}^{*}_{c}$: the first transforms $\bm{\theta}^{*}_{c}$ to an unconstrained space; the second exploits a Taylor approximation to obtain a closed form for the full conditional.
3 Special Case I: Multivariate Distributed Lag Non-Linear Model
Suppose for each unit $i$ we observe each of $P$ exposures at $L$ times; that is, $\mathbf{x}_{ip}=[x_{ip1},\dots,x_{ipL}]^{T}$ for $p=1,\dots,P$. We write the full vector of exposures $\mathbf{x}_{i}^{*}=\left[\mathbf{x}_{i1}^{T},\dots,\mathbf{x}_{iP}^{T}\right]^{T}$ of length MM=PLPL. The goal is to quantify possibly non-linear associations between the $K$ outcomes and the $P$ time-varying exposures. We set the number of indices to be equal to the number of distinct exposures $J=P$ and choose $A_{kj}=\left[\mathbf{0}_{L\times(j-1)L}~|~\mathbf{I}_{L\times L}~|~\mathbf{0}_{L\times(P-j)L}\right],$ the $L\times PL$ horizontal block matrix with $j^{th}$ block equal to the identity matrix, and all other entries zero. This yields a multivariate distributed lag non-linear model (MV-DLNM):
$y_{ik}$ $=f_{k1}({\mathbf{x}}_{i1}^{T}{\bm{\omega}}_{k1})+\cdots+f_{kP}({\mathbf{x}}_{iP}^{T}{\bm{\omega}}_{kP})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},\quad k=1,\dots,K.$ (2)
Here $f_{kp}(\cdot)$ quantifies the potentially non-linear association between the $k^{th}$ outcome and the $p^{th}$ time-varying exposure $\mathbf{x}_{ip}$, and ${\bm{\omega}}_{kp}=({\omega}_{kp1},\dots,{\omega}_{kpL})^{T}$ quantifies the relative strength of that exposure’s association at each of the $L$ times. In particular, adopting the cluster-inducing priors of Section 2.2 here borrows strength across exposure-outcome pairs to exploit similarity in both exposure-response relationships and temporal profiles.
Smoothness over time: Exposure measurements over time may be highly multicollinear, making estimation of $\bm{\omega}_{kp}$ challenging. One solution is to encourage smoothness over time—i.e., encourage $\bm{\omega}_{kpl}$ and $\bm{\omega}_{kpl^{\prime}}$ to be similar when $|l^{\prime}-l|$ is small—corresponding to our prior belief that effects should not change drastically over small time periods. A Gaussian prior with precision matrix $\lambda^{\omega}\mathbf{D}_{s}^{T}\mathbf{D}_{s}$ (where $\mathbf{D}_{s}$ is a known $L-s]\times L$ difference matrix of order $s$) shrinks differences between coefficients at similar lags, with the degree of shrinkage controlled by $\lambda^{\omega}$, a parameter to be estimated (gasparrini2017penalized). However this leaves the magnitude of ${\bm{\omega}}_{kp}$ unconstrained. To address this, we propose an analogous Fisher-Bingham (kent1982fisher) prior $\bm{\omega}_{kj}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{D}_{s}^{T}\mathbf{D}_{s})$, which corresponds to the same Gaussian constrained to the unit $L$-sphere. To further satisfy the sign constraint, we can restrict this prior to the half hypersphere where $\omega_{kpL}\geq 0$.
To facilitate sampling, we can optionally incorporate the dimension reduction approach approximating ${\bm{\omega}}_{kp}\approx\mathbf{\Psi}{\bm{\theta}}_{kp}$, yielding the proposed prior $\bm{\theta}_{kp}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{\Psi}^{T}\mathbf{D}_{s}^{T}\mathbf{D}_{s}\mathbf{\Psi})$, and we can again impose a sign constraint ${{\theta}}_{kpm}>0$, restricting to the half-sphere. We adopt this as the centering distribution $G_{0,\theta}$ for the cluster-inducing priors.
4 Special Case II: Adaptive Multiple Index Model
Suppose we observe each of $P$ exposures at one time only, and let $\mathbf{x}_{i}^{*}=[x_{i1},\dots,x_{iP}]^{T}$ be the $P$-vector of exposures measured once each ($M=P$). Setting $J=1$ and $\mathbf{A}_{1}=\mathbf{I}_{P\times P}$ yields a (multivariate) single index model (SIM):
@@ 4 Special Case II: Adaptive Multiple Index Model
for $k=1,\dots,K$. This SIM can be restrictive, requiring that the associations between each exposure and the $k^{th}$ outcome have the same shape but differ in magnitude. Further adopting the cluster-inducing priors then encourages borrowing strength across the $K$ outcomes.
As a flexible alternative to SIMs, multiple index models (MIMs) for environmental mixtures partition the exposure vector ${\mathbf{x}}_{i}$ (mcgee2021bayesian), but this requires knowing the true index structure a priori. Instead, our proposed framework naturally allows for unknown index structures. First set $J>1$ to be the maximum order of the index, then let $\mathbf{A}_{j}=\mathbf{I}_{P\times P}$ for $j=1,\dots,J$. The (multivariate) adaptive multiple index model of maximum order $J$ is:
$y_{ik}$ $=\sum_{j=1}^{J}f_{kj}({\mathbf{x}}_{i}^{T}{\bm{\omega}}_{kj})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~~k=1,\dots,K.$ (3)
In contrast to model (2), the summation is over multipollutant indices that share common exposures $\mathbf{x}_{i}$. Applying the priors proposed in Section 2.2 now encourages clustering across outcomes and multipollutant indices rather than different exposures. This has two key consequences. First, it again improves efficiency by borrowing strength across outcomes. Second, clustering across indices facilitates collapsing to a lower order (<J<J) index model. For example if JJ=2, then when $Z^{\theta}_{k1}=Z^{\theta}_{k2}$ and $Z^{\beta}_{k1}=Z^{\beta}_{k2}$, model (3) reduces to a single index model—without assuming this a priori. This marks a major advantage over standard MIMs (like in mcgee2021bayesian), as we need only specify the maximum order $J$, rather than the true index structure (i.e., the number of indices and their composition).
@@ 4.1 Exposure Importance Measures
For some choices of $\mathbf{A}_{j}$, weight vectors $\bm{\omega}_{kj}$ have intuitive interpretations: in the MV-DLNM they quantify the strength of association at different lags; in a SIM, they quantify the strength of association between mixture components and outcome. In a MIM with unknown index structure as in Section 4, however, they do not have a natural interpretation, and it is difficult to infer which exposures are driving the mixture association. To facilitate interpretation, we propose alternative exposure importance measures that can be used for complex mixture models like the proposed MIM.
Rewriting $f^{*}_{k}(\mathbf{x})=\sum_{j=1}^{J}f_{kj}({\mathbf{x}}^{T}{\bm{\omega}}_{kj})$ in (3) and omitting the subscript $i$, we define
$\Phi_{kp}=1-\frac{Var_{\mathbf{x}_{-p}}\bigg\{E_{x_{p}\mid\mathbf{x}_{-p}}\Big[f^{*}_{k}(x_{p},\mathbf{x}_{-p})\Big]\bigg\}}{Var[f^{*}_{k}(\mathbf{x})]},$ (4)
@@
where $\mathbf{x}_{-p}$ represents all but the $p^{th}$ exposure. The denominator represents the variability explained by the mixture, and $\Phi_{kp}$ represents the proportion of this variability explained only by the $p^{th}$ exposure. Naturally, $\Phi_{kp}$ will be small if $x_{p}$ is not predictive of the outcome. It will also be small if $x_{p}$ is highly collinear with $\mathbf{x}_{-p}$, as its association can be easily captured by other components. $\Phi_{kp}$ thus quantifies the importance of the $p^{th}$ exposure to the mixture effect on the $k^{th}$ outcome. In Appendix F we generalize this to arbitrary groups of exposures.
@@
Computing these quantities is straightforward given posterior samples of $f_{k}^{*}(\mathbf{x})$. Marginal variances are calculated using the empirical distribution of the exposures. Estimating $\Phi_{kp}$ requires integrating over the conditional distribution of $x_{p}$ given the other exposures. If the dimension of $\mathbf{x}$ is small, we can estimate the conditional mean in (4) non-parametrically:
$\widehat{E}_{x_{p}\mid\mathbf{x}_{-p}}\{f^{*}_{k}(x_{p},\mathbf{x}_{-p})\}=\frac{\sum_{i=1}^{n}f^{*}_{k}(x_{ip},\mathbf{x}_{(-p)})K(\mathbf{x}_{i(-p)}-\mathbf{x}_{-p})}{\sum_{i=1}^{n}K(\mathbf{x}_{i(-p)}-\mathbf{x}_{-p})}$
@@
where $K(\cdot)$ is an appropriate kernel. Alternatively, when $\mathbf{x}$ is high-dimensional, one could make parametric assumptions, such as assuming that $x_{p}$ given $\mathbf{x}_{-p}$ is normally distributed, and the conditional mean can be estimated using any regression approach.
@@ 5 Simulation studies
We demonstrate the effectiveness of the proposed methods in two different simulation studies. In Simulation A, we consider time-varying exposures and demonstrate the benefit of borrowing strength across exposures and outcomes in DLNMs. In B, we consider cross-sectional exposure measurements and try to capture complex, non-linear relationships between a mixture of exposures using MIMs. In both, we average results across 250 simulated datasets.
@@ 5.1 Simulation A: Distributed lag models with multiple exposures
We first focus on settings where exposures are measured repeatedly over time as in Section 3. Throughout, we set the number of outcomes to be $K=4$, the number of exposures to be $P=5$, and the number of time points to be $L=52$. We varied the sample size $n=\{500,1500\}$. We generated exposures from a vector autoregressive model in order to induce autocorrelation within each exposure as well as dependence among different exposures. We assume exposures have correlation 0.85 across successive time points and correlation $0.6^{|j-j^{\prime}|}$ between exposures $j$ and $j^{\prime}$ at each time point (see Appendix G.1 for full details).
@@
Responses were generated from a multivariate normal distribution with mean:
$E[\bm{Y}\mid\bm{X}]$ $=\left[\begin{array}[]{l}f_{1}\left(\mathbf{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\mathbf{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\mathbf{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{2}\left(\mathbf{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\mathbf{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{2}\left(\mathbf{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\mathbf{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{1}\left(\mathbf{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{1}\left(\mathbf{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\mathbf{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{3}\left(\mathbf{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\mathbf{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{4}\left(\mathbf{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{4}\left(\mathbf{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\mathbf{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{4}\left(\mathbf{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\mathbf{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{3}\left(\mathbf{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{3}\left(\mathbf{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\mathbf{x}_{5}^{T}\bm{\omega}_{1}\right)\end{array}\right],$
@@
where $f_{1}(a)=0.04a$, $f_{2}(a)=0.24(0.3a)^{2}$, $f_{3}(a)=0.09a$, $f_{4}(a)=2\sin(0.2a)$; $\bm{\omega}_{1}$ is flat, $\bm{\omega}_{2}$ is decreasing, and $\bm{\omega}_{3}$ is increasing. The lag profiles are the same across outcomes, and the exposure response functions are shared across some, but not all of the exposures.
[…]
@@
Fig. 1 displays average pairwise clustering probabilities. Overall the proposed framework was able to capture similarity in both $f_{kp}(\cdot)$ and $\bm{\omega}_{kp}$. There was a high degree of clustering for the $\bm{\beta}_{kp}$ parameters across exposure-outcome pairs with the same exposure-response function, and a low degree of clustering elsewhere. One exception was that there was a moderate degree of clustering between exposure-outcome pairs with $f_{1}(\cdot)$ as the true exposure response function and those with $f_{3}(\cdot)$. This is due to the fact that both had the same shape of exposure response relationship (both linear), despite being of different strengths. This improves substantially with higher sample size where erroneous clustering occurs much less frequently (bottom left; Fig. 1). Pairwise clustering probabilities were somewhat lower for $\bm{\omega}_{kp}$ (top right and bottom right panel; Fig. 1), as the lag weights were only weakly identified due to high autocorrelation (and exacerbated by correlation among different exposures).